# Machine-Learned Potentials for Solvation Modeling

### v2

@@ Abstract
Solvent environments play a central role in determining molecular structure, energetics, reactivity, and interfacial phenomena. However, modeling solvation from first principles remains difficult due to the complex interplay of interactions and unfavorable computational scaling of first-principles treatment with system size. Machine-learned potentials (MLPs) have recently emerged as efficient surrogates for quantum chemistry methods, offering first-principles accuracy at greatly reduced computational cost. MLPs approximate the underlying potential energy surface, enabling efficient computation of energies and forces in solvated systems, and are capable of accounting for effects such as hydrogen bonding, long-range polarization, and conformational changes. This review surveys the development and application of MLPs in solvation modeling. We summarize the theoretical basis of MLP-based energy and force predictions and present a classification of MLPs based on training targets, model types, and design choices related to architectures, descriptors, and training protocols. Integration into established solvation workflows is discussed, with case studies spanning small molecules, interfaces, and reactive systems. We conclude by outlining open challenges and future directions toward transferable, robust, and physically grounded MLPs for solvation-aware atomistic modeling.
Machine-Learned Potentials (MLP), Machine-Learned Atomistic Potentials (MLAP), Machine-Learned Force-Fields (MLFF), Machine-Learned Interatomic Potentials (MLIP), Solvation Modeling, Hybrid Solvation, Microsolvation
@@ I   Introduction
Recently, machine learning (ML) based strategies referred as machine-learned atomistic potentials (MLAPs), machine-learned interatomic potentials (MLIPs), machine-learned force-fields (MLFF) or briefly, machine-learned potentials (MLPs) have emerged as a promising alternative to traditional first-principles modeling. Such approaches aim to model the potential energy surface (PES) of a system with first-principles accuracy, while enabling scalable atomistic simulations.
@@
Figure 1: Distribution of review articles cited in the present work, categorized by thematic focus (labeled using BibTeX-style citation keys, lastnameYYYYfirstword): traditional solvation modeling based on ab initio and force-field methods (blue), general ML strategies for molecular and materials modeling (green), and ML reviews with explicit discussions on solvation applications (orange). As of yet, there is no focused review solely on “MLPs for solvation modeling”, three reviewsWan, He, and Shi (2024); Yang et al. (2024); Keith et al. (2021) that include dedicated sections on solvation modeling are highlighted in orange. This classification highlights the emerging emphasis on MLPs in the context of solvation modeling.
@@ I.1   What is solvation and why it matters?
The aforementioned examples raise several foundational questions: What are the driving forces behind solvent-mediated molecular processes? When do solute–solvent interactions dominate, and when are solvent–solvent correlations equally significant as solute-solvent interactions? Can such effects be categorized using thermodynamics alone, or do structural motifs also play a decisive role? These are not new questions, but represent some of the problems that are being actively addressed by using experimental and computational techniques Ando et al. (2021); Foumthuim and Giacometti (2023); Jaworek et al. (2021); Buntkowsky, Vogel, and Winter (2018); Ahanger and Dar (2024); Devereux et al. (2020); Jaganade et al. (2020); Raghunathan, Jaganade, and Priyakumar (2020); Raghunathan et al. (2018); Devereux et al. (2014).
@@ I.2   Traditional solvation modeling paradigms
In explicit solvent models, each solvent molecule is modeled individually, enabling a detailed description of both solute–solvent and solvent–solvent interactions. These models are commonly implemented using MD or Monte Carlo (MC) simulations to model the bulk solvation environment, where the solute is embedded in a solvent box comprising a few hundred to thousands of solvent molecules Foumthuim et al. (2020); Foumthuim and Giacometti (2023). When using first-principles methods, a carefully selected unit cell configuration can be treated within the periodic density functional theory (DFT) formalism. While such explicit treatments offer a high degree of realism, they become computationally intractable with the increase in number of solvent molecules. Whereas force fields (FFs) formalism enables simulations of large solvation boxes, the resulting free energy estimates can be highly sensitive to the selected FFs Shirts et al. (2003). For example, simulations of protein stability in osmolyte solutions, such as urea and TMAO, have shown notable variations depending on the FFs employed in MD simulations Canchi and García (2013); Ganguly, van der Vegt, and Shea (2016).
By contrast, implicit solvation models avoid treating individual solvent molecules altogether. Instead, the electrostatic potential provided by the solvent environment is approximated as a continuous polarizable medium Tomasi (2004), thus significantly reducing the computational effort, facilitating rapid and qualitative predictions of solvation-dependent properties and energies. Common frameworks include the polarizable continuum model (PCM) Tomasi, Mennucci, and Cammi (2005), the conductor-like screening model (COSMO) Klamt and Schüürmann (1993), and universal models Cramer and Truhlar (2008) that can describe diverse chemical species (neutral, charged, polar, nonpolar, organics, inorganics) without requiring system-specific parameter tuning such as the solvation model based on density (SMD) Marenich, Cramer, and Truhlar (2009). More physically grounded approaches, such as the reference interaction site model (RISM) Hirata and Rossky (1981); Hirata and Levy (1987) and its three-dimensional extension (3D-RISMSindhikara and Hirata (2013); Imai, Kovalenko, and Hirata (2004), attempt to recover solvent structure through statistical mechanics. In molecular simulations, energy-based methods like molecular mechanics-generalized Born surface area (MM-GBSA) and molecular mechanics-Poisson Boltzmann surface area (MM-PBSA) also fall into the implicit category that can provide fast free energy estimates from snapshots of MD trajectories Genheden and Ryde (2015). Still, these models often miss essential features such as directional hydrogen bonding, ion pair specificity, or the dynamic rearrangement of solvation shells.
@@
where the solute is treated with QM and the MM treatment captures the full cluster minus the solute, hence the solute-solvent interaction is captured at the MM level. In Eq. 1, one can introduce
@@
While the basic ONIOM approach involves three separate energy evaluations per structure, QM/MM methods, especially when based on polarizable or self-consistent electrostatic embedding, may require multiple iterative passes even for single-point energy calculations, due to the coupling between QM and MM regions. Overall, both ONIOM and QM/MM schemes allow for the use of large number of solvent molecules to account for dynamic, explicit solvent interactions while keeping the total computational cost manageable.
@@ I.3   Scope of MLPs for solvation modeling
Solvents play a central role in shaping chemical reactivity at the molecular level, primarily through electrostatic and non-covalent interactions such as hydrogen bonding, dipole–dipole forces, and van der Waals (vdW) interactions. These effects can selectively stabilize different species along the reaction coordinate—reactants, intermediates, or products—alter activation barriers, and ultimately influence both the rate and yield of reactions. Recent advances have shown that MLPs can capture these solvent-mediated effects with near-quantum accuracy and significantly reduced computational cost, enabling realistic modeling of solvation environments in complex chemical processes Anmol and Karmakar (2024); Zhang, Juraskova, and Duarte (2024); Gastegger, Schütt, and Müller (2021); Chen, Zhang, and Zhang (2023).
@@
To understand these complex phenomena at the molecular and electronic scale, one must accurately model the underlying PES. An accurate framework for simulating solvent effects with quantum mechanical accuracy, capturing features like electronic polarization, dynamic hydrogen bonding, and charge transfer is ab initio molecular dynamics (AIMD) Marx and Hutter (2000, 2009); Tuckerman et al. (1996); Carloni, Rothlisberger, and Parrinello (2002). However, its computational complexity restricts its application domain to small systems and short timescales. MLPs offer a practical solution by reproducing the PES with near-AIMD accuracy, enabling longer simulations, enhanced sampling of the configuration space, and facilitating the calculation of statistically converged thermodynamic observables. The data-driven nature of MLPs makes them particularly well-suited for capturing the subtle, many-body interactions that arise during the dynamics of solvated systems. When combined with active learning, these models can explore configuration space efficiently, zeroing in the training efforts on chemically relevant regions while maintaining fidelity to the electronic structure.
In recent years, MLPs—especially those built on deep NN s and active learning frameworks—have gained popularity in computational chemistry Behler (2016); Gastegger, Schütt, and Müller (2021); Zhang et al. (2018a); Mishin (2021); Friederich et al. (2021); Duignan (2024). These models offer a scalable way to approximate PES s with near first-principles accuracy at a fraction of the computational cost. Training on energies and forces allows for the construction of PES s that not only accelerate MD simulations but also facilitate applications in spectroscopy, thermodynamics, and chemical reaction dynamics Meuwly (2021). MLPs are also increasingly replacing classical FF s in the MM region of hybrid QM/MM simulations.
@@
At the heart of any MLP lies a descriptor, which is a numerical representation of the atomic environment, and a mathematical model that maps atomistic configurations to scale, vector, and/or tensor-valued molecular properties Thiemann et al. (2024). The performance of a given MLP hinges on the choice of descriptor and modeling strategy, which can vary widely across NN and kernel-based frameworks Pinheiro et al. (2021). Common approaches include symmetry functions Behler and Csányi (2021) and graph-based encodings. Equally important is the training dataset: it must be diverse and representative enough to capture the relevant configurational space. While high-level electronic structure data remains the gold standard, generating such datasets—especially for systems involving transition metals, enzyme active sites, or complex solvent networks—remains computationally intractable.
@@
In this review, we aim to provide a broad yet focused account of how MLPs are advancing solvation modeling. Section II lays the theoretical foundation, covering various ingredients of MLPs such as training objectives, energy and force formulations, many-body expansions, and the role of $\Delta$-machine learning in improving accuracy. In Section III, we present a taxonomy of MLP architectures—including neural network, kernel-based, linear, and gradient-domain models—and examine how their design choices relate to the specific challenges of solvation. We also explore how MLPs are incorporated into practical modeling workflows, from training strategies to software deployment. Section IV highlights selected case studies that showcase the use of MLPs to construct PES of solvated species, and to directly predict solvent-modulated properties. Further, we highlight the essential role of sampling strategies in the pipeline of solvation-aware MLP modeling. Section V turns to key open challenges, including limitations in model transferability, gaps in benchmarking practices, and the inherent complexity of real solvation environments. Finally, Section VI offers a forward-looking perspective on how the field might evolve toward unified and transferable MLP frameworks for solvation modeling.
@@ II   Physical Foundations of MLPs
MLPs are reshaping the landscape of atomistic simulations by offering a practical way to approximate quantum mechanical PESs with remarkable accuracy and far lower computational cost. Traditional quantum chemistry methods—like DFT and wavefunction-based approaches—deliver high fidelity but become prohibitively expensive as system size grows, limiting their use in large and/or long-timescale MD simulations. Whereas, classical FFs are computationally efficient but often fall short in accuracy and transferability, especially for systems that are chemically diverse or reactive.
MLPs offer a compelling middle ground. By learning the PES directly from ab initio reference data, they achieve near-DFT accuracy while retaining the speed of empirical models. This combination has made it possible to carry out atomistic MD simulations of complex systems—ranging from solvated biomolecules to condensed-phase materials—at spatial and temporal scales that are prohibitively expensive for conventional modeling.
At their core, MLPs are model trained on ab initio data on both energies and forces across a wide range of atomic environments. The rest of this section lays out the theoretical underpinnings of these models, including how they are trained, how loss functions are formulated, the architectural strategies they employ, and establish their connection to classical PES fitting and many-body expansions.
@@ II.1   From PES fitting to MLPs
Like the polynomial expansion in Eq. 6, the NN-potential form in Eq. 8 also decomposes the PES into a sum of products. This structural similarity makes both representations suitable for efficient evaluation of quantum observables by reducing high-dimensional integrals into products of one-dimensional ones, which can accelerate methods such as the multi configuration time-dependent Hartree (MCTDH) approach Manzhos and Carrington Jr (2020).
In both approaches to PES fitting, the coefficients, $c_{f}$ in Eq. 6 and $c_{q}^{\prime}$ in Eq. 8, are obtained by minimizing a loss function that measures the squared-deviation between the fitted and reference (e.g., ab initio) energies over a set of $N_{t}$ training configurations
@@
While total energy is the most commonly used global property in supervised PES learning, other scalar molecular properties, $P\left(\{{\bf r}_{j}\};\{Z_{j}\}\right)$—such as dipole moments, HOMO–LUMO gaps, or solvation free energies—can also be learned using the same framework. In such cases, the loss function takes the general form
@@
where $P_{n}^{\text{ML}}$ and $P_{n}^{\text{ref}}$ denote the predicted and reference values of a given property for configuration $n$. Although this approach has been applied successfully in various molecular property prediction tasks Ramakrishnan and von Lilienfeld (2015), it lacks a clear theoretical justification in many cases. Unlike potential energy, which is a well-defined function of nuclear coordinates via the Born–Oppenheimer approximation, there is often no fundamental reason to expect a smooth or unique mapping from structure to global properties such as redox potential or reactivity—particularly when solvent effects, electronic correlation, or conformational variability play a significant role.
@@
where the weights $w_{E}$ and $w_{F}$ control the trade-off between energy and force accuracy during training. In some cases, additional loss terms are included, such as partial charges, dipole moments, or L2 regularization—particularly in charge-aware or kernel-based models Christensen and Von Lilienfeld (2020). Typically, $w_{F}\propto w_{E}/(3N)$, as for every energy label, the loss function includes $3N$ force labels.
In materials modeling, it is also common to include stress tensor contributions in the training loss to better capture elastic and mechanical responses Ocampo et al. (2024). This combined-training paradigm mirrors practices in computational spectroscopy, where accurate PES construction typically involves fitting to high-level ab initio data—including energies, gradients, and, in some cases, even Hessians Carbonniere et al. (2003, 2004).
II.2   Force prediction

### v3

@@ Abstract
Solvent environments play a central role in determining molecular structure, energetics, reactivity, and interfacial phenomena. However, modeling solvation from first principles remains difficult due to the complex interplay of interactions and unfavorable computational scaling of first-principles treatment with system size. Machine-learned potentials (MLPs) have recently emerged as efficient surrogates for quantum chemistry methods, offering first-principles accuracy at greatly reduced computational cost. MLPs approximate the underlying potential energy surface, enabling efficient computation of energies and forces in solvated systems. They are also capable of accounting for effects such as hydrogen bonding, long-range polarization, and conformational changes. This review surveys the development and application of MLPs in solvation modeling. We summarize the theoretical basis of MLP-based energy and force predictions and present a classification of MLPs based on training targets, model types, and design choices related to architectures, descriptors, and training protocols. Integration into established solvation workflows is discussed, with case studies spanning small molecules, interfaces, and reactive systems. We conclude by outlining open challenges and future directions toward transferable, robust, and physically grounded MLPs for solvation-aware atomistic modeling.
Machine-Learned Potentials (MLP), Machine-Learned Atomistic Potentials (MLAP), Machine-Learned Force Fields (MLFF), Machine-Learned Interatomic Potentials (MLIP), Solvation Modeling, Hybrid Solvation, Microsolvation
@@ I  Introduction
Recently, machine learning (ML) based strategies referred as machine-learned atomistic potentials (MLAPs), machine-learned interatomic potentials (MLIPs), machine-learned force fields (MLFF) or briefly, machine-learned potentials (MLPs) have emerged as a promising alternative to traditional first-principles modeling. Such approaches aim to model the potential energy surface (PES) of a system with first-principles accuracy, while enabling scalable atomistic simulations.
@@
Figure 1: Distribution of review articles cited in the present work, categorized by thematic focus (labeled using BibTeX-style citation keys, lastnameYYYYfirstword): traditional solvation modeling based on ab initio and force-field methods (blue), general ML strategies for molecular and materials modeling (green), and ML reviews with explicit discussions on solvation applications (orange). As of yet, there is no focused review solely on “MLPs for solvation modeling”. Three reviews Wan, He, and Shi (2024); Yang et al. (2024); Keith et al. (2021) that include dedicated sections on solvation modeling are highlighted in orange. This classification highlights the emerging emphasis on MLPs in the context of solvation modeling.
@@ I.1  What is solvation and why it matters?
The examples mentioned above raise several foundational questions: What are the driving forces behind solvent-mediated molecular processes? When do solute–solvent interactions dominate, and when are solvent–solvent correlations equally significant as solute-solvent interactions? Can such effects be categorized using thermodynamics alone, or do structural motifs also play a decisive role? These are not new questions, but represent some of the problems that are being actively addressed by using experimental and computational techniques Ando et al. (2021); Foumthuim and Giacometti (2023); Jaworek et al. (2021); Buntkowsky, Vogel, and Winter (2018); Ahanger and Dar (2024); Devereux et al. (2020); Jaganade et al. (2020); Raghunathan, Jaganade, and Priyakumar (2020); Raghunathan et al. (2018); Devereux et al. (2014).
@@ I.2  Traditional solvation modeling paradigms
In explicit solvent models, each solvent molecule is modeled individually, enabling a detailed description of both solute–solvent and solvent–solvent interactions. These models are commonly implemented using MD or Monte Carlo (MC) simulations to model the bulk solvation environment, where the solute is embedded in a solvent box comprising a few hundred to thousands of solvent molecules Foumthuim et al. (2020); Foumthuim and Giacometti (2023). When using first-principles methods, a carefully selected unit cell configuration can be treated within the periodic density functional theory (DFT) formalism. While such explicit treatments offer a high degree of realism, they become computationally intractable with the increase in the number of solvent molecules. Whereas force fields (FFs) formalism enables simulations of large solvation boxes, the resulting free energy estimates can be highly sensitive to the selected FFs Shirts et al. (2003). For example, simulations of protein stability in osmolyte solutions, such as urea and TMAO, have shown notable variations depending on the FFs employed in MD simulations Canchi and García (2013); Ganguly, van der Vegt, and Shea (2016).
By contrast, implicit solvation models avoid treating individual solvent molecules altogether. Instead, the electrostatic potential provided by the solvent environment is approximated as a continuous polarizable medium Tomasi (2004), thus significantly reducing the computational effort, facilitating rapid and qualitative predictions of solvation-dependent properties and energies. Common frameworks include the polarizable continuum model (PCM) Tomasi, Mennucci, and Cammi (2005), the conductor-like screening model (COSMO) Klamt and Schüürmann (1993), and universal models Cramer and Truhlar (2008) that can describe diverse chemical species (neutral, charged, polar, nonpolar, organics, inorganics) without requiring system-specific parameter tuning such as the solvation model based on density (SMD) Marenich, Cramer, and Truhlar (2009). More physically grounded approaches, such as the reference interaction site model (RISM) Hirata and Rossky (1981); Hirata and Levy (1987) and its three-dimensional extension (3D-RISM) Sindhikara and Hirata (2013); Imai, Kovalenko, and Hirata (2004), attempt to recover solvent structure through statistical mechanics. In molecular simulations, energy-based methods like molecular mechanics-generalized Born surface area (MM-GBSA) and molecular mechanics-Poisson Boltzmann surface area (MM-PBSA) also fall into the implicit category that can provide fast free energy estimates from snapshots of MD trajectories Genheden and Ryde (2015). Still, these models often miss essential features such as directional hydrogen bonding, ion pair specificity, or the dynamic rearrangement of solvation shells.
@@
where the solute is treated with QM, and the MM treatment captures the full cluster minus the solute, hence the solute-solvent interaction is captured at the MM level. In Eq. 1, one can introduce
@@
While the basic ONIOM approach involves three separate energy evaluations per structure, QM/MM methods, especially when based on polarizable or self-consistent electrostatic embedding, may require multiple iterative passes even for single-point energy calculations, due to the coupling between QM and MM regions. Overall, both ONIOM and QM/MM schemes allow for the use of a large number of solvent molecules to account for dynamic, explicit solvent interactions while keeping the total computational cost manageable.
@@ I.3  Scope of MLPs for solvation modeling
Solvents play a central role in shaping chemical reactivity at the molecular level, primarily through electrostatic and non-covalent interactions such as hydrogen bonding, dipole–dipole forces, and van der Waals (vdW) interactions. These effects can selectively stabilize different species along the reaction coordinate, reactants, intermediates, or products, alter activation barriers, and ultimately influence both the rate and yield of reactions. Recent advances have shown that MLPs can capture these solvent-mediated effects with near-quantum accuracy and significantly reduced computational cost, enabling realistic modeling of solvation environments in complex chemical processes Anmol and Karmakar (2024); Zhang, Juraskova, and Duarte (2024); Gastegger, Schütt, and Müller (2021); Chen, Zhang, and Zhang (2023).
@@
To understand these complex phenomena at the molecular and electronic scale, one must accurately model the underlying PES. An accurate framework for simulating solvent effects with quantum mechanical accuracy, capturing features like electronic polarization, dynamic hydrogen bonding, and charge transfer, is ab initio molecular dynamics (AIMD) Marx and Hutter (2000, 2009); Tuckerman et al. (1996); Carloni, Rothlisberger, and Parrinello (2002). However, its computational complexity restricts its application domain to small systems and short timescales. MLPs offer a practical solution by reproducing the PES with near-AIMD accuracy, enabling longer simulations, enhanced sampling of the configuration space, and facilitating the calculation of statistically converged thermodynamic observables. The data-driven nature of MLPs makes them particularly well-suited for capturing the subtle, many-body interactions that arise during the dynamics of solvated systems. When combined with active learning, these models can efficiently explore configuration space, focusing training efforts on chemically relevant regions while maintaining fidelity to the electronic structure.
In recent years, MLPs, especially those built on deep NN s and active learning frameworks, have gained popularity in computational chemistry Behler (2016); Gastegger, Schütt, and Müller (2021); Zhang et al. (2018a); Mishin (2021); Friederich et al. (2021); Duignan (2024). These models offer a scalable way to approximate PES s with near first-principles accuracy at a fraction of the computational cost. Training on energies and forces allows for the construction of PES s that not only accelerate MD simulations but also facilitate applications in spectroscopy, thermodynamics, and chemical reaction dynamics Meuwly (2021). MLPs are also increasingly replacing classical FF s in the MM region of hybrid QM/MM simulations. Beyond non-reactive systems, MLPs have also been developed to capture reactive processes, with models such as ReaxFF enabling the description of bond breaking and formation within complex chemical environments Yang et al. (2024).
@@
At the heart of any MLP lies a descriptor, which is a numerical representation of the atomic environment, and a mathematical model that maps atomistic configurations to scale, vector, and/or tensor-valued molecular properties Thiemann et al. (2024). The performance of a given MLP hinges on the choice of descriptor and modeling strategy, which can vary widely across NN and kernel-based frameworks Pinheiro et al. (2021). Common approaches include symmetry functions Behler and Csányi (2021) and graph-based encodings. Equally important is the training dataset: it must be diverse and representative enough to capture the relevant configurational space. While high-level electronic structure data remains the gold standard, generating such datasets, especially for systems involving transition metals, enzyme active sites, or complex solvent networks, remains computationally intractable.
@@
In this review, we aim to provide a broad yet focused account of how MLPs are advancing solvation modeling. Section II lays the theoretical foundation, covering various ingredients of MLPs such as training objectives, energy and force formulations, many-body expansions, and the role of $\Delta$-machine learning in improving accuracy. In Section III, we present a taxonomy of MLP architectures that span neural network, kernel-based, linear, and gradient-domain models, and examine how their design choices relate to the specific challenges of solvation. We also explore how MLPs are incorporated into practical modeling workflows, from training strategies to software deployment. Section IV highlights selected case studies that showcase the use of MLPs to construct PES of solvated species, and to directly predict solvent-modulated properties. Further, we highlight the essential role of sampling strategies in the pipeline of solvation-aware MLP modeling. Section V turns to key open challenges, including limitations in model transferability, gaps in benchmarking practices, and the inherent complexity of real solvation environments. Finally, Section VI offers a forward-looking perspective on how the field might evolve toward unified and transferable MLP frameworks for solvation modeling.
@@ II  Physical Foundations of MLPs
MLPs are reshaping the landscape of atomistic simulations by offering a practical way to approximate quantum mechanical PESs with remarkable accuracy and far lower computational cost. Traditional quantum chemistry methods, such as DFT and wavefunction-based approaches, deliver high fidelity but become prohibitively expensive as system size grows, limiting their use in large and/or long-timescale MD simulations. Whereas classical FFs are computationally efficient but often fall short in accuracy and transferability, especially for systems that are chemically diverse or reactive.
MLPs offer a compelling middle ground. By learning the PES directly from ab initio reference data, they achieve near-DFT accuracy while retaining the speed of empirical models. This combination has made it possible to carry out atomistic MD simulations of complex systems, ranging from solvated biomolecules to condensed-phase materials, at spatial and temporal scales that are prohibitively expensive for conventional modeling.
At their core, MLPs are model trained on ab initio data on both energies and forces across a wide range of atomic environments. In materials modeling, it is also common to include stress tensor contributions in the training loss to better capture elastic and mechanical responses Ocampo et al. (2024). This combined-training paradigm mirrors practices in computational spectroscopy, where accurate PES construction typically involves fitting to high-level ab initio data, including energies, gradients, and, in some cases, even Hessians Carbonniere et al. (2003, 2004).
The rest of this section lays out the theoretical underpinnings of these models, including their training, loss function formulations, the architectural strategies they employ, and establishes their connection to classical PES fitting and many-body expansions.
@@ II.1  From PES fitting to MLPs
Like the polynomial expansion in Eq. 6, the NN-potential form in Eq. 8 also decomposes the PES into a sum of products. This structural similarity makes both representations suitable for efficient evaluation of quantum observables by reducing high-dimensional integrals into products of one-dimensional ones, which can accelerate methods such as the multi configuration time-dependent Hartree (MCTDH) approach Manzhos and Carrington Jr (2020). This sum-of-products perspective also parallels the design of the atomic cluster expansion (ACE) framework, which employs a systematic product-basis ansatz to represent atomic environments accurately Drautz (2019). In this sense, ACE serves as a physically grounded and interpretable analogue of the NN–based product decomposition.
In both approaches to PES fitting, the coefficients, $c_{f}$ in Eq. 6 and $c_{q}^{\prime}$ in Eq. 8, are obtained by minimizing a loss function that measures the squared deviation between the fitted and reference (e.g., ab initio) energies over a set of $N_{t}$ training configurations
@@
While total energy is the most commonly used global property in supervised PES learning, other scalar molecular properties, $P\left(\{{\bf r}_{j}\};\{Z_{j}\}\right)$, such as dipole moments, HOMO-LUMO gaps, or solvation free energies, can also be learned using the same framework. In such cases, the loss function takes the general form
@@
where $P_{n}^{\text{ML}}$ and $P_{n}^{\text{ref}}$ denote the predicted and reference values of a given property for configuration $n$. Although this approach has been applied successfully in various molecular property prediction tasks Ramakrishnan and von Lilienfeld (2015), it often lacks a clear theoretical justification. Unlike potential energy, which is a well-defined function of nuclear coordinates via the Born–Oppenheimer approximation, there is often no fundamental reason to expect a smooth or unique mapping from structure to global properties such as redox potential or reactivity, particularly when solvent effects, electronic correlation, or conformational variability play a significant role.
@@
where the weights $w_{E}$ and $w_{F}$ control the trade-off between energy and force accuracy during training. In some cases, additional loss terms are included, such as partial charges, dipole moments, or L2 regularization, particularly in charge-aware or kernel-based models Christensen and Von Lilienfeld (2020). Typically, $w_{F}\propto w_{E}/(3N)$, as for every energy label, the loss function includes $3N$ force labels.
II.2  Many-body expansion of energy
II.3  Force prediction: Direct and energy-based